OpenStudy (anonymous):

Find the domain of: g(x)=sqrt of x(x-6)

OpenStudy (anonymous):

\[x \left( x-6 \right)\ge 0\] \[x ^{2}-6x \ge0 \] \[add~ both~ sides~\left( \frac{ -6 }{ 2 } \right)^{2}~i.e.,9\] \[x^{2}-6x+9\ge9\] \[\left( x-3 \right)^{2}\ge 3^{2}\] \[\left| x-3 \right|\ge3\] \[x-3\le-3~or~x \le0\] \[or~x-3\ge3,x \ge6\] so domain is\[(-\infty,0]\cup[6,\infty)\]

OpenStudy (anonymous):

I don't understand where to (-6/2)^2 comes from

OpenStudy (anonymous):

the*

OpenStudy (anonymous):

just to complete the squares. \[\left( a-b \right)^{2}=a ^{2}-2ab+b ^{2},here \left( \frac{ -2b }{2 } \right)^{2}\]

OpenStudy (anonymous):

\[we~ can~ also~ say~ if~ \left( x-a \right)\left( x-b \right)\ge0,then~x~does~\not~lie~\in \left( a,b \right)\] assume a<b

OpenStudy (anonymous):

ok, thank you!